OpenStudy (ac3):

Find the volume of the tetrahedron in the first octant bounded by the coordinate planes and the plane passing through (1,0,0),(0,2,0),(0,0,3)

OpenStudy (irishboy123):

steps: find the equation of that plane first. and write it as z = f(x,y). then integrate over that region

OpenStudy (ac3):

alright let me see if I can get this equation.

OpenStudy (irishboy123):

i make it 6x + 3y + 2z = 6 you should check that though or we will be chasing our tails forever!!

OpenStudy (ac3):

lol I only have the final answer for this

OpenStudy (ac3):

I think I can see how you got that.

OpenStudy (ac3):

When I first did this problem I thought it was simply.... \[\int\limits_{0}^{1}\int\limits_{0}^{2}\int\limits_{0}^{3}dzdydx\]

OpenStudy (irishboy123):

no that is for a cube same problem.....we need to the limits of integration right......

OpenStudy (ac3):

so this region is pretty much a rectangular prism

OpenStudy (irishboy123):

here we have a triangle again, on the xy plane 

OpenStudy (irishboy123):



OpenStudy (irishboy123):

the shaded area is our region. we are integrating \(z = f(x,y) =\frac{1}{2}( 6 - 6x - 3y)\) over that region

OpenStudy (ac3):

man i'm having a really hard time getting to visualize these